### v1

@@ 5.2 Ablation Study
Ablation on Realistic Texts and Synthetic Texts. We conduct ablation experiments on the impact of varying quantities of realistic and synthetic texts using the CLIP-B/32 model. As shown in Tab. 6, increasing the quantity of realistic text from 1 to 3 progressively improves model performance, attributable to enhanced text augmentation that provides extensive real-world knowledge. However, expanding from 3 to 5 slightly reduces performance due to information saturation and noise introduction. Meanwhile, increasing the number of synthetic text from 1 to 5 introduces more noise, gradually decreasing performance. Importantly, supervising image training solely with retrieved realistic texts significantly enhances performance, achieving a 71.2% accuracy compared to 69.8% with the LAION15M dataset. This highlights the crucial role of real-world knowledge in advancing vision-language representation learning.
@@
Figure 9: Visualization of the raw interleaved document, the retrieved realistic text, and synthetic text. Image semantic-related information is highlighted in green.
@@
Case Study. In Fig. 9, we present the visualization of retrieved realistic text and synthetic text obtained from an interleaved document using our proposed transformation paradigm. Both realistic and synthetic texts contain extensive descriptive information consistent with the image semantics, such as “lyell glacier”, “crags”, and “valley”. We provide more visualizations in the supplementary material.
6 Conclusion
@@
This paper explores two open-ended questions: 1) How to utilize multimodal interleaved documents for vision-language representation learning. 2) How to effectively leverage both realistic and synthetic texts to enhance representation performance. To this end, we first establish a Real-World Data Extraction pipeline to extract high-quality images and texts. Then we design a hierarchical retrieval method to efficiently associate each image with multiple semantically relevant texts. To enhance fine-grained image understanding, we propose a visual semantic augmented generation module for synthetic text production. Furthermore, we employ a semantic balance sampling strategy to improve dataset diversity, enabling better learning of long-tail concepts. Based on these innovations, we present RealSyn, a dataset driven by both real and synthetic texts with three sizes: 15M, 30M, and 100M. Comprehensive experimental results show that RealSyn is efficient in vision-language representation learning and exhibits excellent scalability.
7 Limitations
To provide more fine-grained visual information, this paper utilizes vision expert models combined with a Large Language Model (LLM) to generate synthetic text. Given the inference cost and efficiency, the utilization of Modified Large Language Models (MLLM) for synthetic text generation is left for the community to explore. Furthermore, the transformation paradigm proposed in this paper is directly applicable to other multimodal document datasets, including MMC4 Zhu et al. (2024) and OmniCorpus Li et al. (2025). We hope our work provides insights into vision-language representation learning.
8 Ethics Statement
We abide by the ACL Code of Ethics. The data resources used in this study are publicly available.
@@ Appendix A Detail Experiment Settings
A.1 Experiment Settings
In Tab. 7, we present the detailed settings used in training CLIP.
@@ A.1 Experiment Settings
Table 7: Hyperparameters used for CLIP pre-training.
A.2 Detail Instruction Prompt
@@
"Please merge the information from the given raw text and the synthetic caption with the help of the highly relevant detection tags. The raw caption offers detailed real-world information, yet it suffers from flaws in sentence structure and grammar. The synthetic caption exhibits impeccable sentence structure but often lacks in-depth real-world details and may contain false information. The highly relevant detection tags are provided to enrich the semantic information of the raw caption, while some are redundant and noisy. You are a great information integration and summary expert, you are also good at enriching semantic information. Ensure a well-structured sentence while retaining the detailed real-world information provided in the raw caption. Avoid simply concatenating the sentences and avoid adding external information to describe. Correctness and simplify sentences finally. Raw caption:<raw caption>, synthetic caption:<synthetic caption>, and highly relevant detection tags:<detection tags>".
@@ Appendix B Detail External Results
Table 8: List of linear probe datasets with the data distribution and evaluation metrics.
@@
Table 9: Linear probe on 20 downstream datasets. Pre-training different scale CLIP models on RealSyn30M and LAION30M, achieves 1.3%-3.0% average performance improvement.
@@
Table 10: Zero-shot transfer on 20 downstream datasets. Pre-training different scale CLIP models on RealSyn30M and LAION30M, achieves 3.5%-5.5% average performance improvement.
@@
Table 11: Zero-shot robustness comparison. Pre-training different scale CLIP models on RealSyn30M and LAION30M, achieves 4.2%-8.6% average performance improvement.
B.1 Downstream Datasets
To comprehensively demonstrate the performance of CLIP trained on RealSyn, we compared the linear probe results of CLIP trained on RealSyn, YFCC Thomee et al. (2016), and LAION Schuhmann et al. (2021) across 20 datasets. These datasets include Food101 Bossard et al. (2014), CIFAR10 Krizhevsky et al. (2009), CIFAR100 Krizhevsky et al. (2009), Birdsnap Berg et al. (2014), SUN397 Xiao et al. (2010), Stanford Cars Krause et al. (2013), FGVC Aircraft Maji et al. (2013), DTD Cimpoi et al. (2014), Pets Parkhi et al. (2012), Caltech101 Fei-Fei et al. (2004), Flowers102 Nilsback and Zisserman (2008), SLT10 Coates et al. (2011), EuroSAT Helber et al. (2019), RESISC45 Cheng et al. (2017), KITTI Geiger et al. (2012), Country211 Radford et al. (2021), UCF101 Soomro et al. (2012), Hateful Memes Kiela et al. (2020), SST2 Radford et al. (2021), and ImageNet Deng et al. (2009). Details on each dataset and the corresponding evaluation metrics are provided in Tab. 8.
B.2 Detailed Model Scaling Results
Linear Probe. In Tab. 9, we present the detailed linear probe results of different scale CLIP model trained on the 30M dataset. The ViT-L/14 trained on RealSyn30M achieves an average performance improvement of 3.0% across 20 datasets compared to the model trained on the LAION30M.
Zero-shot Transfer. As shown in Tab. 10, ViT-L/14 trained on our proposed RealSyn30M outperforms LAION30M on 18 of 20 downstream datasets and achieves an average improvement of 5.5%.
Zero-shot Robustness. We present the detailed zero-shot robustness performance in Tab. 11, compared with LAION30M, the model trained on RealSyn30M boosts average robustness performance by 8.6% on ViT-L/14.
Figure 10: Data Scaling Analysis. Pre-training ViT-B/32 on RealSyn in different data scales.
B.3 Data Scaling Law
[…]
@@ B.3 Data Scaling Law
Table 12: Linear probe on 12 downstream datasets. Pre-training CLIP (ResNet50) on image-text pairs achieves 2.1% performance improvement.
Dataset #Images #Avg Tokens /// Image #Avg Texts /// Image Text Type Source Type
CC12M 12 000 0001200000012\,000\,00012 000 000 -- 1 Realistic Website
YFCC15M 15 000 0001500000015\,000\,00015 000 000 16 1 Realistic Website
CapsFusion 120 000 000120000000120\,000\,000120 000 000 -- 1 Synthetic Image-Text Pair
LAION400M 400 000 000400000000400\,000\,000400 000 000 27 1 Realistic Website
RealSyn15M 15 239 4981523949815\,239\,49815 239 498 40 4 Realistic & Synthetic Interleaved Image-Text
RealSyn30M 30 328 8523032885230\,328\,85230 328 852 38 4 Realistic & Synthetic Interleaved Image-Text
RealSyn100M 100 862 786100862786100\,862\,786100 862 786 36 4 Realistic & Synthetic Interleaved Image-Text
Table 13: Current Dataset Comparison. Comparison with large-scale image-text pre-training datasets.
$T^{1}_{r}$ $T^{2}_{r}$ $T^{3}_{r}$ $T^{1}_{s}$ Linear probe Avg Transfer Avg Robustness Avg
70.3 42.4 25.7
71.2 46.8 30.7
70.2 39.7 24.0
71.4 47.9 31.5
Table 14: Ablation experiment results using different types of text on RealSyn15M. $T^{k}_{r}$: the $k$-th retrieved semantic relevant realistic text. $T^{k}_{s}$: the image semantic augmented synthetic text for $T^{k}_{r}$.
B.4 Ablation on Retrieved Realistic and Synthetic Texts
Tab. 14 shows the ablation experiment results on the text augmentation using different text types. After introducing the image semantic augmented synthetic text, due to its ability to supplement fine-grain visual semantic information, there are 0.2%, 1.1%, 0.8% performance enhancement on the linear probe, zero-shot transfer, and zero-shot robustness compared to solely utilizing the retrieved realistic texts.
B.5 Detailed Results on Pure Image
To further extend our method for pure images, we conduct experiments on ImageNet Deng et al. (2009). For each image, we retrieve three semantically relevant real-world sentences from our pre-constructed sentence database and generate a single semantically augmented synthetic caption based on the top retrieved text. Following SimCLR Chen et al. (2020), we utilize 4096 batch size and pre-train ResNet50 He et al. (2016) for 90 epochs supervised by the text randomly selected from the three retrieved realistic texts and one synthetic text.
As shown in Tab. 12, compared with SimCLR Chen et al. (2020) under the same conditions, the model trained on our constructed image-text pairs shows an average performance improvement of 2.1% across 12 downstream datasets. The results demonstrate that our method can effectively transform pure images into high-quality image-text pairs through retrieval and generation for vision-language pre-training.
@@ Appendix C Further Analysis of RealSyn
C.1 Compare with Existing Datasets
In Tab. 13, we compare our proposed RealSyn dataset with existing widely used large-scale image-text pre-training datasets. Compared to previous datasets, our proposed RealSyn dataset provides four textual descriptions for each image, with an average token length of 36-40, significantly higher than LAION400M and YFCC15M. Furthermore, unlike previous datasets, the RealSyn dataset is sourced from real-world interleaved image-text documents and includes both realistic and synthetic texts, thereby expanding the scope for future research exploration.
C.2 Visualization of Examples
Figure 11: Visualization of image-text pairs in our proposed RealSyn dataset. $T^{k}_{r}$: the $k$-th retrieved realistic text. $T^{k}_{s}$: the image semantic augmented synthetic text for $T^{k}_{r}$. Image semantic-related information is highlighted in green.
Figure 12: Visualization of image-text pairs in our proposed RealSyn dataset. $T^{k}_{r}$: the $k$-th retrieved realistic text. $T^{k}_{s}$: the image semantic augmented synthetic text for $T^{k}_{r}$. Image semantic-related information is highlighted in green.
In Fig. 11 and Fig. 12, we visualize additional image-text pairs randomly selected from our proposed RealSyn dataset. $T^{k}_{r}$ is the $k$-th retrieved semantically relevant real-world sentence and $T^{k}_{s}$ is the semantic augmentic caption for $T^{k}_{r}$. We also highlighted the image semantic-related information in green and marked the image size below the image.

### v3

@@ 5.2. Ablation Study
Ablation on Realistic Texts and Synthetic Texts. We conduct ablation studies to assess the impact of varying quantities of realistic and synthetic texts on the performance of the CLIP-B/32 model. As shown in Table 6, incrementally increasing the amount of realistic text from one to three enhances model performance, attributed to improved text augmentation that integrates extensive real-world knowledge. However, further increasing this quantity from three to five slightly diminishes performance due to information saturation and the introduction of noise. Conversely, augmenting the number of synthetic texts from one to five incrementally degrades performance, reflecting increased noise introduction. Notably, training with only realistic texts significantly boosts performance, achieving a 71.2% accuracy compared to 69.8% with the LAION15M dataset, underscoring the vital role of real-world knowledge in advancing vision-language representation learning.
Ablation on the Combination of the Realistic Texts and Synthetic Texts. Table 6 presents the results of ablation experiments on text augmentation using different text types. Introducing image semantic augmented synthetic text, which supplements fine-grained visual semantic information, leads to performance enhancements of 0.2%, 1.1%, and 0.8% in linear probe accuracy, zero-shot transfer, and zero-shot robustness, respectively, compared to using only retrieved realistic texts.
@@
Figure 10. Visualization of the raw interleaved document, the retrieved realistic text, and synthetic text. Image semantic-related information is highlighted in green.
@@
Case Study. In Figure 10, we present the visualization of retrieved realistic text and synthetic text obtained from an interleaved image-text document using our proposed transformation paradigm. Both realistic and synthetic texts contain extensive descriptive information consistent with the image semantics, such as “lyell glacier”, “crags”, and “valley”. We provide more visualizations in the supplementary material.
6. Conclusion
@@
This paper explores two open-ended questions: 1) How to utilize multimodal interleaved documents for CLIP training. 2) How to effectively leverage both realistic and synthetic texts to enhance CLIP performance. To this end, we first establish a Real-World Data Extraction pipeline to extract high-quality images and texts. Then we design a hierarchical retrieval method to efficiently associate each image with multiple semantically relevant texts. To further enhance fine-grained visual information, we propose an image semantic augmented generation module for synthetic text production. Furthermore, we employ a semantic balance sampling strategy to improve dataset diversity, enabling better learning of long-tail concepts. Based on these innovations, we present RealSyn, a dataset driven by both real and synthetic texts with three sizes: 15M, 30M, and 100M. We compare our dataset with other widely used datasets of equivalent scale for CLIP training. Models pre-trained on RealSyn consistently achieve state-of-the-art performance across various downstream tasks. Furthermore, extensive experiments confirm that RealSyn significantly enhances contrastive vision-language representation learning and demonstrates robust scalability. We hope our work provides insights into vision-language representation learning.
@@ Appendix A Detail Experiment Settings
A.1. Experiment Settings
In Table 7, we present the detailed settings used in training CLIP.
Table 7. Hyperparameters used for CLIP pre-training.
@@ A.1. Experiment Settings
A.2. Detail Instruction Prompt
@@
”Please merge the information from the given raw text and the synthetic caption with the help of the highly relevant detection tags. The raw caption offers detailed real-world information, yet it suffers from flaws in sentence structure and grammar. The synthetic caption exhibits impeccable sentence structure but often lacks in-depth real-world details and may contain false information. The highly relevant detection tags are provided to enrich the semantic information of the raw caption, while some are redundant and noisy. You are a great information integration and summary expert, you are also good at enriching semantic information. Ensure a well-structured sentence while retaining the detailed real-world information provided in the raw caption. Avoid simply concatenating the sentences and avoid adding external information to describe. Correct and simplify sentences finally. Raw caption:<raw caption>, synthetic caption:<synthetic caption>, and highly relevant detection tags:<detection tags>”.
@@ Appendix B Detail External Results
B.1. Downstream Datasets
To demonstrate the performance of CLIP trained on RealSyn, we compared the linear probe results of CLIP trained on RealSyn, YFCC (Thomee et al., 2016), and LAION (Schuhmann et al., 2021) across 20 datasets. These datasets include Food101 (Bossard et al., 2014), CIFAR10 (Krizhevsky et al., 2009), CIFAR100 (Krizhevsky et al., 2009), Birdsnap (Berg et al., 2014), SUN397 (Xiao et al., 2010), Stanford Cars (Krause et al., 2013), FGVC Aircraft (Maji et al., 2013), DTD (Cimpoi et al., 2014), Pets (Parkhi et al., 2012), Caltech101 (Fei-Fei et al., 2004), Flowers102 (Nilsback and Zisserman, 2008), SLT10 (Coates et al., 2011), EuroSAT (Helber et al., 2019), RESISC45 (Cheng et al., 2017), KITTI (Geiger et al., 2012), Country211 (Radford et al., 2021), UCF101 (Soomro et al., 2012), Hateful Memes (Kiela et al., 2020), SST2 (Radford et al., 2021), and ImageNet (Deng et al., 2009). Details on each dataset and the corresponding evaluation metrics are provided in Table 8.
Table 8. List of linear probe datasets with the data distribution and evaluation metrics.
@@ B.1. Downstream Datasets
B.2. Detailed Model Scaling Results
Table 9. Linear probe on 20 downstream datasets. Pre-training different scale CLIP models on RealSyn30M and LAION30M, achieves 1.3%-3.0% average performance improvement.
@@ B.2. Detailed Model Scaling Results
Linear Probe. In Table 9, we present the detailed linear probe results of different scale CLIP models trained on the 30M dataset. The ViT-L/14 trained on RealSyn30M achieves an average performance improvement of 3.0% across 20 datasets compared to the model trained on the LAION30M.
Table 10. Zero-shot transfer on 20 downstream datasets. Pre-training different scale CLIP models on RealSyn30M and LAION30M, achieves 3.5%-5.5% average performance improvement.
@@
Zero-shot Transfer. As shown in Table 10, ViT-L/14 trained on our proposed RealSyn30M outperforms LAION30M on 18 of 20 downstream datasets and achieves an average improvement of 5.5%.
Table 11. Zero-shot robustness comparison. Pre-training different scale CLIP models on RealSyn30M and LAION30M, achieves 4.2%-8.6% average performance improvement.
@@
Zero-shot Robustness. We present the detailed zero-shot robustness performance in Table 11, compared with LAION30M, the model trained on RealSyn30M boosts average robustness performance by 8.6% on ViT-L/14.
Table 12. Comparison with DreamLIP’s rewriting dataset.
Model Data Scale Dataset Linear probe Avg Transfer Avg
ViT-B/16 15M DreamLIP 73.7 43.3
RealSyn 74.9 51.6
30M DreamLIP 77.2 53.1
RealSyn 77.5 54.9
B.3. Comparison with the Rewrite Dataset.
Due to LaCLIP (Fan et al., 2024) and Capsfusion (Yu et al., 2024) do not provide the model weights trained on the same scale dataset as ours, We compared the same scale model ViT-B/16 pre-trained on the RealSyn with DreamLIP (Zheng et al., 2024b).
As shown in Table 12, ViT-B/16 pre-trained on RealSyn15M can achieve 1.3% and 8.3% average improvement compared to DreamLIP proposed 15M Rewriting datasets based on YFCC15M (Thomee et al., 2016). Furthermore, compared with DreamLIP’s proposed 30M, which combines YFCC15M, CC12M (Changpinyo et al., 2021) and CC3M (Changpinyo et al., 2021) dataset, ViT-B/16 training on RealSyn30M can also outperform it in the linear probe and zero transfer. This significant improvement indicates that realistic knowledge is important for the CLIP model.
B.4. Combine RealSyn with LAION
As shown in Table 13, integrating RealSyn15M with LAION15M yields average improvements of 0.6% in linear probe, 0.2% in zero-shot transfer across 20 datasets, and a robustness increase of 0.6% compared to LAION30M. Comparing RealSyn30M with LAION30M, the former shows significant performance gains, with average improvements of 1.3% in linear probe, 3.5% in zero-shot transfer, and a robustness enhancement of 4.2%. The experimental results demonstrate that RealSyn  when combined with existing datasets, achieves significant performance improvements, while also validating the robustness and extensibility of RealSyn.
Table 13. Data Scaling Comparison. Merged 30M: LAION 15M + RealSyn15M.
Model Dataset Linear probe Avg Transfer Avg Robustness Avg
ViT-B/16 LAION15M 69.8 42.7 27.2
RealSyn15M 71.4 47.9 31.5
LAION30M 72.6 48.6 33.6
Merged30M 73.2 48.8 34.2
RealSyn30M 73.9 52.1 37.8
Table 14. Linear probe on 12 downstream datasets. Pre-training CLIP (ResNet50) on image-text pairs achieves 2.1% performance improvement.
@@ B.4. Combine RealSyn with LAION
Table 15. Current Dataset Comparison. Comparison with large-scale image-text pre-training datasets.
Dataset #Images #Avg Tokens // Image #Avg Texts // Image Text Type Source Type
CC12M 12 000 00012\,000\,000 - 1 Realistic Website
YFCC15M 15 000 00015\,000\,000 16 1 Realistic Website
CapsFusion 120 000 000120\,000\,000 - 1 Synthetic Image-Text Pair
LAION400M 400 000 000400\,000\,000 27 1 Realistic Website
RealSyn15M 15 239 49815\,239\,498 40 4 Realistic & Synthetic Interleaved Image-Text
RealSyn30M 30 328 85230\,328\,852 38 4 Realistic & Synthetic Interleaved Image-Text
RealSyn100M 100 862 786100\,862\,786 36 4 Realistic & Synthetic Interleaved Image-Text
B.5. Detailed Results on Pure Image
To further extend our method for pure images, we conduct experiments on ImageNet (Deng et al., 2009). For each image, we retrieve three semantically relevant real-world sentences from our pre-constructed sentence database and generate a single semantically augmented synthetic caption based on the top retrieved text. Following SimCLR (Chen et al., 2020), we utilize 4096 batch size and pre-train ResNet50 (He et al., 2016) for 90 epochs supervised by the text randomly selected from the three retrieved realistic texts and one synthetic text.
As shown in Table 14, compared with SimCLR (Chen et al., 2020) under the same conditions, the model trained on our constructed image-text pairs shows an average performance improvement of 2.1% across 12 downstream datasets. The results demonstrate that our method can effectively transform pure images into high-quality image-text pairs through retrieval and generation for vision-language pre-training.
@@ Appendix C Further Analysis of RealSyn
C.1. Compare with Existing Datasets
In Table 15, we compare our proposed RealSyn dataset with existing widely used large-scale image-text pre-training datasets. Compared to previous datasets, our proposed RealSyn dataset provides four textual descriptions for each image, with an average token length of 36-40, significantly higher than LAION400M and YFCC15M. Furthermore, unlike previous datasets, the RealSyn dataset is sourced from real-world interleaved image-text documents and includes both realistic and synthetic texts, thereby expanding the scope for future research exploration.
C.2. Visualization of Examples
Figure 11. Visualization of image-text pairs in our proposed RealSyn dataset. $T^{k}_{r}$: the $k$-th retrieved realistic text. $T^{k}_{s}$: the image semantic augmented synthetic text for $T^{k}_{r}$. Image semantic-related information is highlighted in green.
Figure 12. Visualization of image-text pairs in our proposed RealSyn dataset. $T^{k}_{r}$: the $k$-th retrieved realistic text. $T^{k}_{s}$: the image semantic augmented synthetic text for $T^{k}_{r}$. Image semantic-related information is highlighted in green.
In Figure 11 and Figure 12, we visualize additional image-text pairs randomly selected from our proposed RealSyn dataset. $T^{k}_{r}$ is the $k$-th retrieved semantically relevant real-world sentence and $T^{k}_{s}$ is the semantic augmentic caption for $T^{k}_{r}$. We also highlighted the image semantic-related information in green and marked the image size below the image.
Appendix D Limitations
To provide more fine-grained visual information, this study employs vision expert models in conjunction with a Large Language Model (LLM) to generate synthetic text. Considering inference costs and efficiency, further exploration of Modified Large Language Models (MLLM) for synthetic text creation is suggested for the community. Additionally, constrained by computational resources, this paper constructs a 100M-scale RealSyn dataset exclusively using OBELICS (Laurençon et al., 2024). Notably, the transformation paradigm presented here is directly applicable to other multimodal document datasets, including MMC4 (Zhu et al., 2024b) and OmniCorpus (Li et al., 2025).